the square root of 18y+9 =y+5
Found 2 solutions by stanbon, Prithwis:
Answer by stanbon(48546)

Square both sides to get:
18y+9= y^2+6x+9
y^2-12y=0
y(y-12)=0
y=0 or y=12
Cheers,
Stan H.

Answer by Prithwis(166)

Square both sides:
18y+9= y^2+10y+25
=> y^2-8y+16=0
=> (y-4)^2 = 0
=> y = 4